"How much water must be evaporated for 120kg of solution which is 3% salt to make a solution of 5% salt?"
Please correct me if I'm wrong:
Let x = amount of water to be evaporated
0.03(120) - 0(x) = 0.05(120-x)
3.6 = 6 - 0.05x
0.05x = 6 - 3.6
x = 2.4 / 0.05
x = 48L
Thank you

Answer by edjones(8007)

.03*120=3.6
3.6/x = .05
x=3.6/.05 multiply each side by x/.05
x=72
120-72=48 L
Right on!
.
Ed